OpenStudy (anonymous):

what is the simplest form of the radical expression? 3^3sqrt2a-6^3sqrt2a

OpenStudy (solomonzelman):

\[3\sqrt[3]{2a}+6\sqrt[3]{2a}\]this?

OpenStudy (anonymous):

yes!

OpenStudy (solomonzelman):

\[3a+6a=9a\]\[3\sqrt[3]{2a}+6\sqrt[3]{2a}=~?\]

OpenStudy (anonymous):

is that what is simplified or are you asking me what it equals?

OpenStudy (solomonzelman):

I am asking you how would you simplify it, knowing that\[3x+6x~~~~->~~~~9x\]\[3\sqrt[3]{2a}+6\sqrt[3]{2a}~~~~~~->~~~~~?\]

OpenStudy (solomonzelman):

This is just adding like terms.

OpenStudy (anonymous):

ooh so would i add 2a + 2a and get 4a? since 2a and 2a are like terms

OpenStudy (nikato):

No, when ur adding roots, u like what's inside and only calculate eth e outside, which is usually left of the sign

OpenStudy (nikato):

So in ur case, I'll just add 3+6

OpenStudy (anonymous):

so combine 3^3 and 6^3

OpenStudy (anonymous):

so i get 9^3?

OpenStudy (mertsj):

\[3x+6x=(3+6)(x)=9x\] \[3\sqrt[3]{2a}+6\sqrt[3]{2a}=(3+6)(\sqrt[3]{2a})= ????\]

OpenStudy (nikato):

No, that 3 is part of the radical sign. Instead of finding the square root, which u see most often, that 3 means cube root

OpenStudy (anonymous):

ohhhh so 9^3sqrt2a

OpenStudy (mertsj):

Bingo

OpenStudy (anonymous):

thank you so much!

OpenStudy (mertsj):

yw

OpenStudy (nikato):

Yes, u write it like that, but remember the 9 and 3 don't really have a relation. It does not mean 9 to the 3rd power

OpenStudy (saiyan):

you all missed the -6 in the equation its not +6